An 800 ft ship has a displacement of 35,000 tons, and the




area defined by the waterline is 38,000 ft2




. Will the ship take




more or less draft when steaming from salt water to fresh water?




How much will it settle or rise?

1
Expert's answer
2021-11-29T19:08:39-0500

"\rho _s>\rho_f," it will settle.

"V=Sh,"

"F_A=\rho_1 gV_1=\rho_2gV_2,"

"\rho_1Sh_1=\rho_2Sh_2=m,"

"\Delta h=h_1-h_2=\frac mS(\frac 1{\rho _1}-\frac 1{\rho_2})=26~cm."